OpenStudy (anonymous):

-16x -1y-3/80x -4y-3 simplify

OpenStudy (anonymous):

-16x -1y - (3/80x) -4y -3? or (-16x -1y-3)/(80x -4y-3)

OpenStudy (anonymous):

1st way u have it written

OpenStudy (radar):

If that is the way then just combine terms getting: 42.66x-5y-6

OpenStudy (anonymous):

i am willing to bet this is \[\frac{-16x^{-1}y^{-3}}{80x^{-4}y^{-3}}\] because i saw this problem today already

OpenStudy (radar):

I think you're right, it looks wrong

OpenStudy (anonymous):

cancel the -16 and the 80 leaves -5 in the denominator. the \[y^{-3}\] cancel as well. this leaves \[\frac{x^{-1}}{-5x^{-4}}\] now bring up the \[x^{-4}\] into the numerator to get \[\frac{x^{-1}x^4}{-5}\] ad the exponents to get \[-\frac{x^3}{5}\]

OpenStudy (anonymous):

yeah i saw this once already

OpenStudy (radar):

Hopefully bradley0516 will verify.

OpenStudy (anonymous):

yes thats the way it was written but no one had responded so I reposted it again

OpenStudy (radar):

Looks like you got an answer now.